OpenStudy (anonymous):

16a-36 -------- 4 Need help!!

OpenStudy (anonymous):

please type the question correctly or you can draw it

OpenStudy (anonymous):

you can split this into 2 fractions and reduce \[\frac{ 16a-36 }{ 4 }=\frac{ 16a }{ 4 }-\frac{ 36 }{ 4 }\]